If consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200. Y is domestic income and Y is private disposable income. Determine the value of the economy's multiplayer which is applicable to government spending and interpret it.


1
Expert's answer
2021-05-11T15:12:27-0400

"Yd = Y - T"

"Y = C + I + G + X - M"

"Y = 100 + 0.75Yd + 200 + 150 + 200 - 50 - 0.25 Y"

"Y = 600 + 0.75(Y - T) - 0.25Y"

"1.25Y = 600 + 0.75(Y - 50 - 0.5Y)"

"1.25Y = 600 - 37.5 + 0.375Y"

"1.25Y - 0.375Y = 562.5"

"0.875Y = 562.5"

"Y = 642.85"

Now suppose government spending increased to 200. 

"Y = 100 + 0.75Yd + 200 + 200 + 200 - 50 - 0.25Y."

"0.875Y = 612.5"

"Y =\frac{ 612.5}{0.875}"

"Y = 700"

If government spending increases by 50 then the Y have increased by

"(700-642.85) = 57.15"

Government spending multiplier "= \frac{∆Y}{∆G}"  


Government spending multiplier "=\frac{57.15}{50}=1.143"

An increase in one government spending will increase 1.43 of income.